HUBBLE SPACE TELESCOPE DAILY REPORT #2878 PERIOD COVERED: 0000Z (UTC) 05/22/01 - 0000Z (UTC) 05/23/01 Daily Status Report as of 143/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 8828 (Cycle 9 Supplemental Darks pt3/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.2 Completed WF/PC-2 9157 (Fundamental Properties Of L-Type Dwarfs In Binaries) The WF/PC-2 was used to characterize the physical properties of eight L-dwarfs in four binary systems. The goal is to obtain astrometric, photometric and spectroscopic measurements of each component that will yield basic information on their atmospheric and dynamical properties. The proposal completed nominally. 1.3 Completed Two Sets of WF/PC-2 9045 (The Relationship Between Radio Luminosity and Radio-Loud AGN Host Galaxy Properties) The Wf/PC-2 was used to determine the relationship between the properties of the host galaxies of radio-loud AGN and their radio luminosities. Previous studies in this area with the HST have concentrated on the 3C sample which shows a tight correlation between luminosity and redshift, such that evolutionary effects cannot be distinguished from those depending upon radio luminosity. Our sample of 46 radio galaxies at z ~ 0.5 comes from four complete, low-frequency-selected samples of radio sources with differing flux limits. Thus the total sample spans an unprecedented three orders of magnitude in radio luminosity at a fixed redshift interval. The proposal completed with no reported problems. 1.4 Completed WF/PC-2 8585 (Cosmological Parameters from Type Ia Supernovae at High Redshift) The WF/PC-2 was used to measure the cosmological parameters, Omega, Lambda, and thus the curvature, Omega_k, using Type Ia supernovae {SNe Ia} as calibrated standard candles. There were no reported problems. 1.5 Completed WF/PC-2 9144 (Calibrating Star Formation: The Impact of Environment) The WF/PC-2 was used to obtain UV, V, and I broad-band and HAlpha, OIII, SII, and HBeta emission line images of three nearby, starburst galaxies to: map the young stellar populations and star clusters, locate the obscuring dust clouds, and diagnose conditions in the ionized ISM. The sample is carefully selected to probe a range of environments for the host galaxy, from isolated to group-interacting, for constant morphology, luminosity and metallicity. The galaxies are all closer than ~4 Mpc, giving ~2 pc resolution with the WF so we can chart the detailed structure of each starburst. This project is designed to: {a} establish whether starbursts are large-scale coherent or chaotic processes by investigating the presence of sequential/propagating star formation; and {b} explore the influence of the external environment on the evolution of the starburst. For this purpose we will measure: {1} the recent star formation history of both stellar clusters and young diffuse population; {2} the dust distribution; and {3} the structures in the shocked and photo-ionized gas. This study will provide a physical basis, in terms of energetics, evolution, and internal structures, for the interpretation of blue galaxies at intermediate and high redshifts. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 5 Successful: 5 Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 13 Successful: 13 2.3 Operations Notes: The STIS instrument remains in safing. A TTR was written for a required re-transmit during a NSSC-1 load uplink at 142/1321Z. Accordingly, using ROP NS-5, the SI C&DH error counter was reset at 142/1332Z. At 142/1554Z, per an operations request, the flight operations team (FOT) entered "P" format and then commanded the STIS primary power relay on (closed PDU2 relay). Upon return to "H" format at 142/1620Z, the FOT was able to verify that the relay telemetry still showed an off (open) state, indicating a failure to pass current through the circuit. At 142/1619Z, per another operations request, the FOT successfully enabled the STIS side 2 zone A heaters. ROP SR-01A was used to clear the SSR EDAC error counter at 142/1850Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: GSFC Internal Simulation #8 (EVA-1 -- -V2 SA wing), a three-day exercise, will begin this morning at 8:00 a.m. Today's session will conclude at approximately 8:00 p.m. Continuation of normal science observations and calibrations. /CAW